OpenStudy (anonymous):

A shopkeeper marks his goods at 40% above cost (the price he pays for the goods) and then offers a 25% discount. If a jacket costs the shopkeeper $200, at what price is it sold? please help i'm clueless ?

OpenStudy (welshfella):

marked price = 200 + 40% of 200 = 200 * 0.4*200 = 200 + 80 = $280

OpenStudy (welshfella):

price after 25% discount = 280 - 0.25*280 can you work that out?

OpenStudy (anonymous):

not really so do i multiply first & then add ? @welshfella

OpenStudy (welshfella):

yes do the 0.25 * 280 first then subtract this from $280

OpenStudy (anonymous):

so its going to be 0.25x280=70 & then 70-280=210 ?

OpenStudy (welshfella):

that's right $210 is your answer

OpenStudy (anonymous):

thank you .

OpenStudy (welshfella):

yw